A large cooler contains the following drinks: 8 lemonades, 6 Sprites, 13 Cokes, and 5 root beers. You randomly pick two cans, one at a time (without replacement). Compute the following probabilities.
(a) What is the probability that you get two cans of Sprite?

(b) What is the probability that you do not get two cans of Coke?

(c) What is the probability that you get either two root beers or two lemonades?

(d) What is the probability that you get one can of Coke and one can of Sprite?

(e) What is the probability that you get two drinks of the same type?

Answer by ewatrrr(24785)

32 Total: 8 lemonades, 6 Sprites, 13 Cokes, and 5 root beers
pick two cans, one at a time (without replacement)
|
The idea is: P = amount of the kind chosen/Total
Remembering on the second pick one less of both: kind and Total
a) P(2-S) = (6/32)(5/31)
b) P(NOT 2 C) = 1 - (13/32)(12/31)
c) P(2 RB 0r 2L) = (5/32)(4/31)+ (8/32)(7/31)
e) P = (6/32)(5/31) + (13/32)(12/31) + (5/32)(4/31)+ (8/32)(7/31)
d) P = (13/32)(6/31) + (6/32)(13/31) (Depends on which kind is picked first)